# Electric field induction in quark-gluon plasma due to thermoelectric effects

### v1

@@ A.2 Case-II (B $\not$= 0)
Hence, one can write the linear independent solutions corresponding to the homogeneous part of differential Eq.  (A.27) in terms of its eigenvectors as
$y_{1}=\begin{pmatrix}-iFe^{\eta_{1}}\\ 0\\ 0\\ e^{\eta_{1}}\end{pmatrix},~{}~{}y_{2}=\begin{pmatrix}0\\ -iFe^{\eta_{2}}\\ e^{\eta_{2}}\\ 0\end{pmatrix},~{}~{}y_{3}=\begin{pmatrix}iFe^{\eta_{3}}\\ 0\\ 0\\ e^{\eta_{3}}\end{pmatrix},~{}~{}y_{4}=\begin{pmatrix}0\\ iFe^{\eta_{4}}\\ e^{\eta_{4}}\\ 0\end{pmatrix}.$
@@
From Eq. (A.30) and Eq. (A.27) we can see that $Y_{p}$ is a column matrix with the same coefficients as that of matrix X. Further, the differentiation of Eq. (A.30) with respect to time gives us $Y_{p}^{{}^{\prime}}$ = $Y^{{}^{\prime}}U+YU^{{}^{\prime}}$, where $Y^{{}^{\prime}}=AY$. Hence, $Y_{p}^{{}^{\prime}}$ = $AY_{p}+YU^{{}^{\prime}}$. Comparison of above equation with Eq. (A.27) tells us
@@
$\alpha_{2}$ $=\frac{-\tau_{R}^{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i})(1+\chi_{i}^{2})}\alpha_{1}$
@@
$\alpha_{4}$ $=\frac{-\tau_{R}^{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i})(1+\chi_{i}^{2})}\alpha_{3}$
$\alpha_{5}$ $=\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{3}$
$\alpha_{6}$ $=-\tau_{R}^{i}\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{3}$
$\alpha_{7}$ $=-\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{B(1+\chi_{i})(1+\chi_{i}^{2})}\alpha_{3}$
$\alpha_{8}$ $=\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{1}$
$\alpha_{9}$ $=-\tau_{R}^{i}\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{1}$
@@
$\delta f_{i}$ $=\frac{-q_{i}\tau_{R}^{i}}{(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}\Big{\{}(1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{\}}(\vec{v_{i}}.\vec{E})\pm\Big{\{}\chi_{i}(1+\chi_{i})$
$+\frac{\chi_{i}^{2}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{\}}\Big{(}\vec{v_{i}}.(\vec{E}\times\hat{b})\Big{)}\Big{]}+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}$
$-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\pm\chi_{i}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{T}\times\hat{b})\Big{)}\mp\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}$
$\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{\dot{T}}\times\hat{b})\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.38)
For simplicity we have considered $\tau_{E}=\tau_{B}$ and $\chi_{i}=\frac{\tau_{R}^{i}}{\tau_{B}}=\frac{\tau_{R}^{i}}{\tau_{E}}$. Here $\pm$ indicates positive sign for positively charged particles(antiparticles) and negative sign for negatively charged particles (antiparticles).
Acknowledgements.
KS acknowledge the doctoral fellowship from UGC, Government of India. JD and RS gratefully acknowledge the DAE-DST, Govt. of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021- IITI (E-37123).

### v2

@@ A.2 Case-II (B $\not$= 0)
Hence, one can write the linearly independent solutions corresponding to the homogeneous part of differential Eq.  (A.27) in terms of its eigenvectors as
$y_{1}$ $=\begin{pmatrix}-iFe^{\eta_{1}}\\ 0\\ 0\\ e^{\eta_{1}}\end{pmatrix},~{}~{}y_{2}=\begin{pmatrix}0\\ -iFe^{\eta_{2}}\\ e^{\eta_{2}}\\ 0\end{pmatrix},$
$y_{3}$ $=\begin{pmatrix}iFe^{\eta_{3}}\\ 0\\ 0\\ e^{\eta_{3}}\end{pmatrix},~{}~{}~{}y_{4}=\begin{pmatrix}0\\ iFe^{\eta_{4}}\\ e^{\eta_{4}}\\ 0\end{pmatrix}.$
@@
From Eq. (A.30) and Eq. (A.27), we can see that $Y_{p}$ is a column matrix with the same coefficients as that of matrix X. Further, the differentiation of Eq. (A.30) with respect to time gives us $Y_{p}^{{}^{\prime}}$ = $Y^{{}^{\prime}}U+YU^{{}^{\prime}}$, where $Y^{{}^{\prime}}=AY$. Hence, $Y_{p}^{{}^{\prime}}$ = $AY_{p}+YU^{{}^{\prime}}$. Comparison of above equation with Eq. (A.27) tells us
@@
$\alpha_{2}$ $=\frac{-\tau_{R}^{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i})(1+\chi_{i}^{2})}\alpha_{1},$
@@
$\alpha_{4}$ $=\frac{-\tau_{R}^{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i})(1+\chi_{i}^{2})}\alpha_{3},$
$\alpha_{5}$ $=\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{3},$
$\alpha_{6}$ $=-\tau_{R}^{i}\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{3},$
$\alpha_{7}$ $=-\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{B(1+\chi_{i})(1+\chi_{i}^{2})}\alpha_{3},$
$\alpha_{8}$ $=\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{1},$
$\alpha_{9}$ $=-\tau_{R}^{i}\frac{\chi_{i}}{B(1+\chi_{i})}\alpha_{1},$
@@
$\delta f_{i}$ $=\frac{-q_{i}\tau_{R}^{i}}{(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}\Big{\{}(1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{\}}$
$(\vec{v_{i}}.\vec{E})\pm\Big{\{}\chi_{i}(1+\chi_{i})+\frac{\chi_{i}^{2}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{\}}\Big{(}\vec{v_{i}}.(\vec{E}\times\hat{b})\Big{)}\Big{]}$
$+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}$
$-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\pm\chi_{i}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{T}\times\hat{b})\Big{)}$
$\mp\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{\dot{T}}\times\hat{b})\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.38)
For simplicity we have considered $\tau_{E}=\tau_{B}$ and $\chi_{i}=\frac{\tau_{R}^{i}}{\tau_{B}}=\frac{\tau_{R}^{i}}{\tau_{E}}$. Here $\pm$ indicates a positive sign for positively charged particles(antiparticles) and a negative sign for negatively charged particles (antiparticles).